Determine whether the function f (x) = x − x1 is odd, even or neither.


1
Expert's answer
2021-11-23T16:55:03-0500

A function is even if "f(x) = f(-x)";

and if "f(-x) = -f(x)", then the function is odd;

if "f(-x) \neq f(x) , f(-x)", then the function is neither



"f(x) = x-x_1\\
f(-x) = -x -x_1 = -(x+x_1)"


Since the function is neither even "(x-x_1)" or odd "-(x-x_1)", then it is neither"."